# Continual Release Moment Estimation with Differential Privacy

### v1

@@ 7 Summary and Discussion
In this work we studied the problem of jointly estimating the first and second moment of a continuous data stream in a differentially private way. We presented the Joint Moment Estimation (JME) method, which solves the problem by exploiting the recently proposed matrix mechanism with carefully tuned noise level. As a result, JME produces unbiased estimates of both moments while requiring less noise to be added than baseline methods, at least is the high privacy regime. We applied JME to private Gaussian density estimation and model training with Adam, demonstrating improved performance in high-privacy regimes both theoretically and practically.
Despite the promising results, several open question remain. In particular, would like to explore if postprocessing is indeed the optimal strategy in a low-privacy regime, or if a better privacy-utility trade-off is still possible. Furthermore, we plan to explore the possibility of problem-specific factorizations, which could be fused with the proposed method.
@@ Acknowledgments
We thank Monika Henzinger for her valuable feedback and insightful discussions on earlier versions of this draft. We are also grateful to Mher Safaryan for his contributions to discussions on DP-Adam. Additionally, we thank Ryan McKenna for suggesting Joint Clipping (CS) as a baseline. Jalaj Upadhyay’s research was funded by the Rutgers Decanal Grant no. 302918 and an unrestricted gift from Google. A part of this work was done while visiting the Institute of Science and Technology Austria (ISTA). Christoph Lampert’s research was partially funded by the Austrian Science Fund (FWF) 10.55776/COE12.
@@ Appendix A Proofs of the Main Theorems
The privacy of Algorithm 1 follows from the properties of matrix mechanism (Li et al., 2015), with a precise estimate of the sensitivity of the joint estimation process that we will introduce and discuss later in this section.
@@
where $\lambda>0$ is an arbitrary trade-off parameter that we will adjust optimally later. To make the estimate private, we privatize the rightmost matrix in (15), which contains the data, by adding suitable scaled Gaussian noise. The subsequent matrix multiplication then acts on private data, so its result is also private.
@@
For this, we denote by $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;\max_{i}\|x_{i}\|_{2}\leq\zeta\wedge x_{i}\in\mathbb{R}^{d}\}$ be the set of input sequences where each $d$-dimensional vectors has bounded $\ell_{2}$ norm. The sensitivity of a computation is the maximal amount by which the result differs between two input sequences $X,X^{\prime}\in\mathcal{X}$, which are identical except for a single data vector at an arbitrary index (denoted by $X\sim X^{\prime}$).
@@
$\text{sens}_{\lambda}^{2}(C_{1},C_{2})$ $=\max_{i=1,\dots,n}\sup_{\|x\|\leq\zeta,\|y\|\leq\zeta}\alpha^{2}_{i}\|x-y\|_{2}^{2}+\lambda\beta^{2}_{i}\|x\otimes x-y\otimes y\|_{2}^{2}$
$\;=\frac{\zeta^{2}}{\alpha^{2}_{1}}\left[\sup_{\|x\|\leq 1,\|y\|\leq 1}\|x-y\|_{2}^{2}+\frac{\lambda\zeta^{2}\beta^{2}_{1}}{\alpha^{2}_{1}}\|x\otimes x-y\otimes y\|_{2}^{2}\right]$
@@
Corollary 1 implies that the joint estimate (15) will be $(\epsilon,\delta)$-private, if we use noise of strength at least, $\sigma=\zeta^{2}\alpha_{1}r_{d}(\lambda\zeta^{2}\beta_{1}/\alpha_{1})\sigma_{\epsilon,\delta}$, where $\sigma_{\epsilon,\delta}$ is the noise strength required for the Gaussian mechanism with sensitivity $1$. This finishes the proof of the Lemma 3.5.
@@
Inserting $B_{1}=A_{1}C_{1}^{-1}$ and $\sigma=2\zeta\sigma_{\epsilon,\delta}\|C_{1}\|_{1\to 2}$, Equation (3) follows. Analogoulsy, for $Z_{2}\sim[\mathcal{N}(0,\sigma^{2}\text{I})]^{d\times d}$,
@@
$r^{\text{diag}}_{d}(\lambda)=\sup\limits_{\|x\|_{2}\leq 1,\|y\|_{2}\leq 1}\!\!\!\big{[}\|x-y\|_{2}^{2}+\lambda\|\operatorname{\operatorname{diag}}(x\otimes x)-\operatorname{\operatorname{diag}}(y\otimes y)\|_{2}^{2}\big{]}=r_{d}(\lambda)$
@@
$\|x-y\|_{2}^{2}+\lambda\|x\circ x-y\circ y\|_{2}^{2}\leq\|x^{\prime}-y^{\prime}\|_{2}^{2}+\lambda\|x^{\prime}\circ x^{\prime}-y^{\prime}\circ y^{\prime}\|_{2}^{2}.$
@@ Appendix B Additional Materials
Definition B.1 (Face-Splitting Product).
@@
  • Frobenious norm $\|A\operatorname{\!\bullet\!}B\|_{\text{F}}=\|B\operatorname{\!\bullet\!}A\|_{\text{F}}$, even so $A\operatorname{\!\bullet\!}B\neq B\operatorname{\!\bullet\!}A$.
Appendix C Technical Proofs
C.1 Bounds on the expected approximation error for PP with non-trivial factorization
Lemma C.1.
@@ C.1 Bounds on the expected approximation error for PP with non-trivial factorization
Let $\beta=\langle x,y\rangle$, where $-1\leq\beta\leq 1$. The expression $-\beta-2\lambda\beta^{2}$ is maximized at $\beta=-\frac{1}{2\lambda}$ when $\lambda\geq\frac{1}{2}$, or at $\beta=-1$ when $\lambda<\frac{1}{2}$. If $\beta=-1$, then $x=-y$ and the objective function becomes $4$. If $\beta=-\frac{1}{2\lambda}$, the objective function is equal to $2+2\lambda+\frac{1}{2\lambda}$, which concludes the proof.
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(X+C_{1}^{-1}Z_{1}))-A_{2}(X\operatorname{\!\bullet\!}X)\|_{\text{F}}^{2}$
$\quad=2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\operatorname{\!\bullet\!}C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}}_{\mathcal{S}_{1}}+2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\langle A_{2}(X\operatorname{\!\bullet\!}C_{1}^{-1}Z_{1}),A_{2}(C_{1}^{-1}Z_{1}\operatorname{\!\bullet\!}X)\rangle_{\text{F}}}_{\mathcal{S}_{2}}+\underbrace{\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}}_{\mathcal{S}_{3}}.$
We compute those terms separately.
@@
$\begin{split}\mathbb{E}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}&=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &\quad+2\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}\\ &\quad+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}\\ &=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (52)
@@
$\begin{split}\mathbb{E}(C_{1}^{-1}Z_{1})_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{2},i}(C_{1}^{-1}Z_{1})_{t_{2},j}&=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}\right)^{2}=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (53)
@@
$\mathcal{S}_{3}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+(d+1)Q_{t_{1},t_{2}}^{2})$
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=d(d+1)\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))+2(d+1)\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
@@
$=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}\sum\limits_{r=1}^{n}(A_{2})_{k,t}(C_{1}^{-1})^{2}_{t,r}=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}(A_{2})_{k,t}Q_{t,t}.$
@@
$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$
@@
$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$
@@
$\mathbb{E}_{Z}\|VZ\operatorname{\!\bullet\!}VZ\|_{\text{F}}^{2}=d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}.$
@@
$\mathbb{E}_{Z}\langle(VZ)\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ)\rangle=\frac{1}{d}\mathbb{E}_{Z}\|VZ_{1}\operatorname{\!\bullet\!}VX\|_{\text{F}}^{2}.$
@@
$\mathbb{E}_{Z}\langle(VZ)\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ)\rangle$ $=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}(VZ_{1})_{k,i}(VX)_{k,j}(VZ_{1})_{k,j}(VX)_{k,i}=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\left(\sum\limits_{i=1}^{d}(VZ_{1})_{k,i}(VX)_{k,i}\right)^{2}$
@@
$\mathbb{E}_{Z}\|AZ\|_{\text{F}}^{2}=\|A\|_{\text{F}}^{2}.$

### v2

@@ 7 Summary and Discussion
In this work we studied the problem of jointly estimating the first and second moment of a continuous data stream in a differentially private way. We presented the Joint Moment Estimation (JME) method, which solves the problem by exploiting the recently proposed matrix mechanism with carefully tuned noise level. As a result, JME produces unbiased estimates of both moments while requiring less noise to be added than baseline methods, at least in the high privacy regime. We applied JME to private Gaussian density estimation and model training with Adam, demonstrating improved performance in high-privacy regimes both theoretically and practically.
Despite the promising results, several open question remain. In particular, we would like to explore if postprocessing is indeed the optimal strategy in a low-privacy regime, or if a better privacy-utility trade-off is still possible. Furthermore, we plan to explore the possibility of problem-specific factorizations, which could be fused with the proposed method.
@@ Acknowledgments
We thank Monika Henzinger for her valuable feedback and insightful discussions on earlier versions of this draft. We are also grateful to Mher Safaryan for his contributions to discussions on DP-Adam. Additionally, we thank Ryan McKenna for suggesting Joint Clipping as a baseline. Jalaj Upadhyay’s research was funded by the Rutgers Decanal Grant no. 302918 and an unrestricted gift from Google. A part of this work was done while visiting the Institute of Science and Technology Austria (ISTA). Christoph Lampert’s research was partially funded by the Austrian Science Fund (FWF) 10.55776/COE12.
@@ Appendix A Proofs of the main theorems
The privacy of Algorithm 1 follows from the properties of matrix mechanism [34], with a precise estimate of the sensitivity of the joint estimation process that we will introduce and discuss later in this section.
@@
where $\lambda>0$ is an arbitrary trade-off parameter that we will adjust optimally later. To make the estimate private, we privatize the rightmost matrix in (15), which contains the data, by adding suitably scaled Gaussian noise. The subsequent matrix multiplication then acts on private data, so its result is also private.
@@
For this, we denote by $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;\max_{i}\|x_{i}\|_{2}\leq\zeta\wedge x_{i}\in\mathbb{R}^{d}\}$ be the set of input sequences where each $d$-dimensional vector has bounded $\ell_{2}$ norm. The sensitivity of a computation is the maximal amount by which the result differs between two input sequences $X,X^{\prime}\in\mathcal{X}$, which are identical except for a single data vector at an arbitrary index (denoted by $X\sim X^{\prime}$).
@@
$\text{sens}_{\lambda}^{2}(C_{1},C_{2})$ $=\max_{i=1,\dots,n}\sup_{\|x\|\leq\zeta,\|y\|\leq\zeta}\alpha^{2}_{i}\|x-y\|_{2}^{2}+\lambda\beta^{2}_{i}\|x\otimes x-y\otimes y\|_{2}^{2}$ (22)
$\;=\zeta^{2}\alpha^{2}_{1}\left[\sup_{\|x\|\leq 1,\|y\|\leq 1}\|x-y\|_{2}^{2}+\frac{\lambda\zeta^{2}\beta^{2}_{1}}{\alpha^{2}_{1}}\|x\otimes x-y\otimes y\|_{2}^{2}\right]$
@@
Corollary A.3 implies that the joint estimate (15) will be $(\epsilon,\delta)$-private, if we use noise of strength at least, $\sigma=\zeta^{2}\alpha_{1}r_{d}(\lambda\zeta^{2}\beta_{1}/\alpha_{1})\sigma_{\epsilon,\delta}$, where $\sigma_{\epsilon,\delta}$ is the noise strength required for the Gaussian mechanism with sensitivity $1$. This finishes the proof of the Lemma 3.5.
@@
Inserting $B_{1}=A_{1}C_{1}^{-1}$ and $\sigma=2\zeta\sigma_{\epsilon,\delta}\|C_{1}\|_{1\to 2}$, Equation (3) follows. Analogously, for $Z_{2}\sim[\mathcal{N}(0,\sigma^{2}\text{I})]^{d\times d}$,
@@
$r^{\text{diag}}_{d}(\lambda)=\sup\limits_{\|x\|_{2}\leq 1,\|y\|_{2}\leq 1}\!\!\!\big{[}\|x-y\|_{2}^{2}+\lambda\|\operatorname{\operatorname{diag}}(x\otimes x)-\operatorname{\operatorname{diag}}(y\otimes y)\|_{2}^{2}\big{]}=r_{d}(\lambda)$ (27)
@@
$\|x-y\|_{2}^{2}+\lambda\|x\circ x-y\circ y\|_{2}^{2}\leq\|x^{\prime}-y^{\prime}\|_{2}^{2}+\lambda\|x^{\prime}\circ x^{\prime}-y^{\prime}\circ y^{\prime}\|_{2}^{2}.$ (29)
@@ Appendix B Additional materials
Table 1: Common workload matrices for moment estimation in (1).
method symbol coefficients
prefix sum $E_{1}$ $\mathbbm{1}{\{i\leq t\}}$
exponential (weight $\beta$) $E_{\beta}$ $\beta^{t-i}\mathbbm{1}\{i\leq t\}$
standard average $V$ $\frac{1}{t}\mathbbm{1}\{i\leq t\}$
sliding window (size $k$) $W_{k}$ $\frac{1}{k}\mathbbm{1}\{t-k\,<\,i\,\leq\,t\}$
(a) trivial factorization
(b) square root factorization
(a) $d=1$, $n=1000$
(b) $d=100$, $n=1000$
@@
Appendix C Algorithms
Algorithm 2 Differentially Private JME Adam
$m_{0}\leftarrow 0$   // first moment initialization.
$v_{0}\leftarrow 0$   // second moment initialization.
for $i=1,2,\dots,n$ do
     $S_{i}\leftarrow\{d_{1},\dots,d_{m}\}\subseteq D$ select a data batch
     $x_{i}\leftarrow\sum_{j=1}^{m}\min(1,\zeta/||g_{j}||)g_{j}$
     $\widehat{x}_{i}\leftarrow x_{i}+\zeta[C_{\beta_{1}}^{-1}Z_{1}]_{[i,\cdot]}$
     $m_{i}\leftarrow m_{i-1}\beta_{1}+(1-\beta_{1})\widehat{x}_{i}$
     $v_{i}\leftarrow v_{i-1}\beta_{2}+(1-\beta_{2})\widehat{x^{2}}_{i}$
     $\widehat{m}_{i}=m_{i}/(1-\beta_{1}^{i})$   // bias-correction
     $\widehat{v}_{i}=v_{i}/(1-\beta_{2}^{i})$   // bias-correction
end for
$\Theta=(\theta_{1},\dots,\theta_{n})$
Algorithm 3 $\lambda$-JME
input stream vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms)  (default: $\text{I}_{n\times n}$)
privacy parameters $(\epsilon,\delta)$
for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
[…]
end for
input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
privacy parameters $(\epsilon,\delta)$
$\sigma\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
privacy trade-off $\alpha\in(0,1)$
$\sigma_{1}\leftarrow\frac{\sigma}{\sqrt{\alpha}}$
$s_{1}\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
$\sigma_{2}\leftarrow\frac{\sigma}{\sqrt{1-\alpha}}$
for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
[…]
end for
Algorithm 5 $\tau$-CS (CS with 2nd moment rescaling parameter $\tau>0$)
input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with
workload matrices $A=(a^{t}_{k})\in\mathbb{R}^{n\times n}$
input dimension $d$, bound on input norm $\zeta>0$
privacy parameters $(\epsilon,\delta)$
(optional) noise shaping matrix $C$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
for $t=1,2,\dots,n$ do
[…]
end for
Algorithm 6 PP
input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
noise shaping matrix $C_{1}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
privacy parameters $(\epsilon,\delta)$
$s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z]_{[t,\cdot]}$
[…]
end for
Appendix D Technical Proofs
Definition D.1 (Face-Splitting Product).
@@ Appendix D Technical Proofs
  • Frobenius norm $\|A\operatorname{\bullet}B\|_{\text{F}}=\|B\operatorname{\bullet}A\|_{\text{F}}$, even though $A\operatorname{\bullet}B\neq B\operatorname{\bullet}A$.
D.1 Bounds on the expected approximation error for PP with non-trivial factorization
Lemma D.2.
@@ D.1 Bounds on the expected approximation error for PP with non-trivial factorization
Let $\beta=\langle x,y\rangle$, where $-1\leq\beta\leq 1$. The expression $-2\beta-2\lambda\beta^{2}$ is maximized at $\beta=-\frac{1}{2\lambda}$ when $\lambda\geq\frac{1}{2}$, or at $\beta=-1$ when $\lambda<\frac{1}{2}$. If $\beta=-1$, then $x=-y$ and the objective function becomes $4$. If $\beta=-\frac{1}{2\lambda}$, the objective function is equal to $2+2\lambda+\frac{1}{2\lambda}$, which concludes the proof.
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\operatorname{\bullet}(X+C_{1}^{-1}Z_{1}))-A_{2}(X\operatorname{\bullet}X)\|_{\text{F}}^{2}$ (52)
$\quad=2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\operatorname{\bullet}C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}}_{\mathcal{S}_{1}}+2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\langle A_{2}(X\operatorname{\bullet}C_{1}^{-1}Z_{1}),A_{2}(C_{1}^{-1}Z_{1}\operatorname{\bullet}X)\rangle_{\text{F}}}_{\mathcal{S}_{2}}$
$\qquad\qquad+\underbrace{\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}}_{\mathcal{S}_{3}}.$
We compute those terms separately. We add component-wise independent Gaussian noise $Z_{1}\sim\mathcal{N}(0,\sigma^{2})$ to the first moment, where $\sigma=2\zeta\sigma_{\epsilon,\delta}$.
@@
$\begin{split}\mathbb{E}&(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}\\ &=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &\quad+2\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}\\ &\quad+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}\\ &=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (56)
@@
$\begin{split}\mathbb{E}(C_{1}^{-1}Z_{1})_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{2},i}(C_{1}^{-1}Z_{1})_{t_{2},j}&=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}\right)^{2}\\ &=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (57)
@@
$\mathcal{S}_{3}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ (58)
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=d(d+1)\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$ (59)
$\quad+2(d+1)\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
@@
$=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}\sum\limits_{r=1}^{n}(A_{2})_{k,t}(C_{1}^{-1})^{2}_{t,r}$
$=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}(A_{2})_{k,t}Q_{t,t}.$
@@
$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$ (61)
@@
$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$ (62)
Thus, we can eliminate the last term ($d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$) in the error sum via bias correction. Substituting back $\sigma=2\zeta\sigma_{\epsilon,\delta}$, we obtain the proposed equality. ∎
@@
$\mathbb{E}_{Z}\|VZ\operatorname{\bullet}VZ\|_{\text{F}}^{2}=d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}.$ (63)
@@
$\mathbb{E}_{Z}\langle(VZ)\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ)\rangle=\frac{1}{d}\mathbb{E}_{Z}\|VZ_{1}\operatorname{\bullet}VX\|_{\text{F}}^{2}.$ (65)
@@
$\mathbb{E}_{Z}\langle(VZ)\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ)\rangle$ $=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}(VZ_{1})_{k,i}(VX)_{k,j}(VZ_{1})_{k,j}(VX)_{k,i}$ (66)
$=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\left(\sum\limits_{i=1}^{d}(VZ_{1})_{k,i}(VX)_{k,i}\right)^{2}$
@@
$\mathbb{E}_{Z}\|AZ\|_{\text{F}}^{2}=\|A\|_{\text{F}}^{2}.$ (67)